OpenStudy (anonymous):

I need help with x^2+y^2=2x I need to get x=? out of this can anyone explain this to me?

OpenStudy (ddcamp):

Have you done completing the square yet?

OpenStudy (anonymous):

here what is the value of veriable y???

OpenStudy (anonymous):

this is hoe far I've gotten x^2 -2x+y^2 =0 (x-1)^2-y^2=0 (x-1)^2=-y^2 but then I'm not sure what to do I'm totally having a brain fart, too much turkey I guess. :) but how do I get it down to x=?

OpenStudy (anonymous):

And this is part of a very long integral but I'm in the process of taking partial derivatives so I just let y ride as a constant for now.

OpenStudy (anonymous):

do u want to solve all of above 3 equations step seperetly?? or these are the part of one question?

OpenStudy (anonymous):

In need to use the x=(?) in something and I cant remember how to get this down to that but I just need to get the expression to put in an integral.

OpenStudy (anonymous):

X^2+Y2=2x is the given function and I need to solve it for x

OpenStudy (anonymous):

X^2 + Y^2 = 2X is it sorry

OpenStudy (anonymous):

i think partial derivative will let u to reach on conclusion

OpenStudy (anonymous):

I'm not trying to take the partial of this I need to take dx of another function and I need the x= of this to use as an upper bound in an integral

OpenStudy (anonymous):

sorry dear i have no idea bout that since m a student of Electrical Engineering,,,i have not yet faced this type of equation...

OpenStudy (anonymous):

oh if your an engineering student you will eventually but this is in cal 3 but I just forgot this procedure. I'm also an engineering student (civil), good luck

OpenStudy (anonymous):

hehehe i know its not just a difficult prob but now i too forgot how to solve this...

OpenStudy (anonymous):

no worries i'll figure it out eventually